# SFFT-based Homogenization: Using Tensor Trains to Enhance FFT-Based Homogenization

### v1

@@ Abstract
Homogenization is a key technique for approximating the macroscopic properties of materials with microscale heterogeneity. The FFT-based Homogenization method has gained widespread usage due to its computational efficiency and accuracy in handling complex microstructures. However, despite its advantages, the method is limited by speed and memory constraints, particularly when applied to high-resolution discretizations. These limitations affect its scalability and efficiency, especially in large-scale simulations or when dealing with highly detailed microstructures. These challenges arise from the fundamental reliance on the Fast Fourier Transform, which imposes inherent restrictions on further advancements.
In this paper, we propose a novel SFFT-based Homogenization algorithm that utilizes a Quantized Tensor Train variant of the Quantum Fourier Transform. This method is tailored to the geometry under consideration and offers significant improvements in time complexity and memory efficiency compared to the traditional FFT-based approach while remaining executable on classical hardware. The method is applicable only if a suitable Quantized Tensor Train representation exists for the stiffness operator associated with the underlying geometry.
@@ 1 Introduction
Homogenization is a mathematical technique used to approximate the macroscopic properties of materials with microscale heterogeneity, such as composites and alloys [1, 2, 3]. Computational Homogenization enables simulations of such materials, even when microstructural features are orders of magnitude smaller than the macroscopic scale [4].
@@
In recent years, several approaches have sought to improve state-of-the-art FFT-based Homogenization methods. One such strategy explores the use of quantum computing, specifically by utilizing the improved time complexity of the Quantum Fourier Transform (QFT) over the traditional FFT [10, 11]. While these approaches are promising, they face significant challenges due to the need for fault-tolerant quantum computing. As a result, these methods are unlikely to be feasible in the near future, particularly within the current Noisy Intermediate-Scale Quantum (NISQ) era.
Another promising approach employs tensor networks (TNs) in a ’quantum-inspired’ manner, aiming to achieve computational speedups through compression. Although these methods do not face the same challenges as quantum computing, they still fall short in overcoming the time complexity barrier imposed by the FFT [12, 13].
@@
In this paper, we present a novel SFFT-based Homogenization algorithm and investigate its applicability to various geometries in both 2D and 3D. A key aspect of our analysis is the memory consumption and speed of our approach, which we compare to the state-of-the-art FFT-based Homogenization method. Our goal is to achieve scalability similar to the QFT while ensuring practical computability with classical computers.
The outline of the remainder of this paper is as follows: Section 2 presents the derivation of the FFT-based Homogenization algorithm and discusses its limitations, while Section 3 introduces the (Quantized) Tensor Train format. Section 4 presents the ’Zip-Up’ algorithm, which converts our QFT quantum circuit into the desired SFFT Tensor Train, and concludes with our novel SFFT-based Homogenization algorithm. In Section 5, we compare the performance of our approach to the state-of-the-art FFT-based Homogenization method, focusing primarily on speed and memory consumption. Before closing, we provide a short summary of the paper – along with the statement of the main conclusions drawn – in Section 6.
@@ 2 FFT-based Homogenization
We continue the derivation by identifying that the local stress tensor must satisfy the force-equilibrium equation
@@
where the Fast Fourier Transform (FFT) $\mathcal{F}[\mathbf{\,\,}]$ is employed for efficient computation. Throughout this paper, variables in momentum space will be denoted by a hat symbol to distinguish them from their real-space counterparts.
@@
Working with tensor representations can often be cumbersome, both in terms of clarity and computational efficiency. To address this, we adopt Voigt notation, which represents stress and strain tensors as 3 (6) dimensional vectors in this notation, the stiffness tensor is represented as a $3\times 3$ ($6\times 6$) matrix acting on them in 2 (3) dimensional space. This reduction is made possible by exploiting the inherent symmetries of the tensors [20].
The effective stiffness tensor $\mathbf{C}^{\mathrm{eff}}$ can conveniently be determined in Voigt notation, where the strain-stress relationship (1) reduces to a simple matrix-vector product. By selectively controlling the mean strain $\mathbf{E}$ where only its $k$-th component to unity while all others are set to zero - the relationship simplifies to its component form
@@
We conclude this section by noting that the computational complexity of the FFT-based algorithm 1 is dominated by the FFT operations, scaling with $\mathcal{O}\left(N^{D}\log N\right)$, where $N$ is the number of discretization points per dimension $D$. Thus, improving the FFT-based Homogenization procedure relies on either enhancing the performance of FFT operations, such as through hardware acceleration, or circumventing the classical FFT entirely.
@@ 3 Tensor Train Methods
Operations Rank Result Time Complexity
$z=x*\mathnormal{const}$ $r(z)=r(x)$ $\mathcal{O}\left(dr(x)\right)$
$z=Ax$ (solve z) $r(z)$ $\mathcal{O}\left(ndr(A)r^{3}(z)\right)$
$Az=x$ (solve z) $r(z)$ $\mathcal{O}\left(ndr(A)r^{3}(z)\right)$
Tensor Networks (TNs) have a rich and diverse history, having been developed independently within both the mathematics and physics communities [22, 23]. This cross-disciplinary development led to a variety of powerful and complementary approaches across the field. In the context of TNs, high-order tensors are decomposed into a network of lower-order tensors, interconnected through specific contraction patterns. This allows for efficient manipulation and representation of complex data.
There are various TN formats, including Tree Tensor Networks (TTNs) and Projected Entangled Pair States (PEPS) [24, 22, 25]. These methods are well-suited for representing and processing high-dimensional tensors, with each having its own advantages in different contexts. However, this work specifically focuses on the Tensor Train (TT) format, along with some of its special cases, for representing high-order tensors $\mathcal{T}$ [26, 27].
The elements of $\mathcal{T}$ are related to its TT decomposition as follows
@@
where each physical index $i_{k}$ runs from 00 to $d_{k}-1$. The novel 3-order tensors $C^{\alpha_{k}}_{\alpha_{k-1}}(i_{k})$ of shape $(r_{k}\times d_{k}\times r_{k-1})$ are regarded as the cores of the TT. The highest dimensionality of all the virtual subspaces is denoted as the rank $r$ of the TT providing insight into the feasibility of employing TTs.
The primary advantage of TNs over full tensor representations lies in their data compression capabilities. An $n$-order tensor, with each index having a dimension of $d$, requires exponential memory $O(d^{n})$ to store its data. In contrast, the same tensor in the TT format scales linearly with its dimension and order $O(ndr^{2})$, while now additionally depending on its rank. Thus, effective compression is achieved only if the rank $r$ remains moderate.
The feasibility of the TT decomposition relies on the underlying structure of the tensor. Unstructured tensors, such as random tensors, do not have a low-rank representation and are therefore unsuitable for TT decomposition. This limitation is not specific to TT but applies generally to TNs. However, different TN formats may represent the same structured data in various ways, resulting in varying levels of compression efficiency [28].
Moreover, TTs are not merely a means of compressed storage. They offer a robust mathematical framework that enables efficient operations, leading to significant computational speedups when applicable. The extent of this acceleration depends primarily on the TT rank. The time complexity and rank evolution of TTs under basic linear algebra are summarized in Table 1. Given the broad mathematical foundation of TTs, we refer readers to I. V. Oseledets introduction instead of discussing basic linear algebra in detail [29].
A generalization called the Tensor Train Operator (TTO), can be obtained by extending the TT formula 9 to include an additional physical index per core
@@
The physical index pairs of each core can be interpreted as incoming and outgoing dimensions, which allows a TT to be contracted with a TTO. Therefore, the TTO acts as a complex transformation on a TT.
A more complex class of operations includes optimization algorithms, often employed to solve linear systems of equations (LSE). In this paper, we discuss an adaptation of the alternating minimal energy (AMEn) algorithm [30]. Rather than using the AMEn algorithm to solve LSEs, we modify it to accelerate matrix-vector product calculations by adjusting the loss function in the optimization process. Our goal is to find an approximate solution $v$ of the equation $Au=\tilde{v}$, with the exact TTs $u$ and $\tilde{v}$ by minimising the following cost function
@@
The general procedure closely follows the derivation outlined in Dolgov’s work [30], with deviations in certain details. Additional information on the adapted method is provided in Appendix B.
The primary advantage of this class of optimization methods over direct TTO-TT contractions becomes evident in the high-rank regime The introduced approach not only reduces time complexity but also results in lower final ranks, often eliminating the need for an additional TT rounding step. As shown in Table 1, this speeds up subsequent operations.
To contract tensors efficiently, it can be beneficial to represent a regular TT as a TTO to perform computations. This is mainly due to the enhanced efficiency of the corresponding optimization methods. Assume we have a TT $u$ as defined by equation (9). We can raise it to a TTO $U$, following the format in equation (3), by employing delta functions
@@
where the novel TTO cores $G^{\alpha_{k}}_{\alpha_{k-1}}(i_{k},j_{k})$ are 4-order tensors of shape $(d_{k}\times r_{k}\times d_{k}\times r_{k-1})$. The TTO derived from relation (11) applied to a TT $v$ produces the same outcome as the Hadamard product of $u$ and $v$. However, the rank of the resulting TT exhibits enhanced scalability, as it is no longer governed by the product of the ranks of $u$ and $v$, see Table 1. A detailed proof is provided in Appendix C.
In general, operations increase the rank when applied on a TT, eventually becoming computationally intractable, as shown in Table 1. To address this issue, additional rounding operations must be employed. Given a tensor $v$ in TT format, the goal is to find an approximation $\tilde{v}$ such that
@@
fulfilling the desired accuracy $\delta_{\mathrm{acc}}$ in terms of the frobenius norm ||$||_{F}$. The process involves sequential orthogonalization of TT cores followed by truncated SVDs, cutting small singular values. To reach the desired accuracy, the truncation threshold during the procedure must satisfy the condition
@@
This iterative process preserves essential information while maintaining a reduced rank and computational tractability [29].
@@ 4 Quantum-Inspired Approach
To align the TT format with quantum circuits, we restrict the dimensionality of the physical indices to $d=2$. This specialization, called the Quantized Tensor Train (QTT) format, directly links the core dimensionality to the single-qubit subspaces in quantum circuits. Thus, an $n$-core QTT represents an $n$-qubit quantum state, with unitary gates acting on these qubits being represented as Quantized Tensor Train Operators (QTTOs).
The mathematical relationship between the QTT format and quantum circuits reveals a closeness in both their mathematical structure and diagrammatic representations. This duality forms the cornerstone of our quantum-inspired approach.
@@ 4.1 The Quantum Fourier Transform
$H=\frac{1}{\sqrt{2}}$ $\begin{pmatrix}1&1\\ 1&-1\\ \end{pmatrix},\hskip 28.45274ptSWAP=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\\ \end{pmatrix},$
@@
The complete circuit is shown in Figure 1. Alternatively, the QFT can be written as a product of two operators $\mathcal{QFT}_{n}=R_{n}\,Q_{n}$. The $Q_{n}$ circuit is generally responsible for the transformation itself, while the $R_{n}$ contribution reorders the output through a succession of $SWAP$ Gates. The latter one is needed to achieve the same ordering convention as is used within its classical counterpart.
The QFT exhibits an exponentially improved time complexity of $\mathcal{O}\left(d\log(N)^{2}\right)$ compared to the regular FFT, which scales as $\mathcal{O}\left(N^{d}\log(N)\right)$, making it a desirable alternative.
Figure 1: Quantum Circuit for the Quantum Fourier Transform $\mathcal{QFT}_{n}=\mathcal{R}_{n}\mathcal{Q}_{n}$ for $n=4$.
The transfer of the QFT into its corresponding QTT formulation has been done previously, however, it was deemed impractical due to its high ranks [15, 31]. Recently, Chen et al. [14] showed that this may be true for the full QFT, but not for its reduced form $\mathcal{Q}_{n}$. Specifically, they showed that when the reordering operator $\mathcal{R}_{n}$ is not considered, the ranks obtained for the QTT formulated reduced QFT decay exponentially with $n$.
The QTTO version of the QFT was dubbed the Superfast Fourier Transform (SFFT) [15], as it was observed that for certain applications, the SFFT outperforms the regular FFT. These cases typically require the underlying data for the SFFT to be structured
We will adopt a similar approach by restricting ourselves to the reduced form and deriving the SFFT. However, this leads to a different convention for the momentum space, which must be considered for the Green-Eshelby operator in the upcoming section. It is worth noting that even on current NISQ-era quantum hardware, the reordering operation can cause issues due to the numerous error-prone SWAP operators [32].

### v3

@@ Abstract
Homogenization is a fundamental technique for estimating the macroscopic properties of materials with microscale heterogeneity. Among Homogenization methods, the FFT-based Homogenization algorithm has become widely used due to its computational efficiency and ability to handle complex microstructures. Nevertheless, even with GPU acceleration, FFT-based Homogenization for industrial applications remains excessively time-consuming, particularly when generating elastic training data for AI models. This is due to the curse of dimensionality, which arises from the algorithms reliance on the Fast Fourier Transform, creating a fundamental bottleneck.
In this paper, we propose a quantum-inspired SFFT-based Homogenization algorithm that leverages the improved time complexity of a Tensor Train variant of the Quantum Fourier Transform. By additionally exploiting structural properties of the underlying microstructure, our method achieves exponential improvements in time complexity and memory efficiency compared to the traditional FFT-based technique - all while remaining executable on classical hardware. We evaluate the performance of our algorithm across increasingly complex microstructures, demonstrating its potential advantages and limitations.
@@ 1 Introduction
Homogenization is a mathematical technique used to approximate the macroscopic properties of materials with microscale heterogeneity, such as composites and alloys [1, 2, 3, 4, 5]. Computational Homogenization enables simulations of such materials, even when microstructural features are orders of magnitude smaller than the macroscopic scale [6, 7, 8, 9, 10].
@@
In recent years, several approaches have sought to improve state-of-the-art FFT-based Homogenization methods. One such strategy uses GPUs to significantly accelerate the procedure, however it is severely limited by the memory constrains of the used GPUs [16]. Another strategy explores the use of quantum computing, specifically by utilizing the improved time complexity of the Quantum Fourier Transform (QFT) over the traditional FFT [17, 18]. While these approaches are promising, they face significant challenges due to the need for fault-tolerant quantum computing. As a result, these methods are unlikely to be feasible in the near future, particularly within the current Noisy Intermediate-Scale Quantum (NISQ) era.
Another promising approach employs Tensor Networks in a quantum-inspired manner, aiming to achieve computational speedups through compression. Although these methods do not face the same challenges as quantum computing, they still fall short in overcoming the time complexity barrier imposed by the FFT [19, 20].
@@
Table 1: List of important symbols and abbreviations used in this work.
Symbol Description
TN Tensor Network
TT Tensor Train
TTO Tensor Train Operator
SFFT Superfast Fourier Transform
NISQ Noisy Intermediate-Scale Quantum
$n$ TT cores / Qubits per Spatial Dimension
$N$ Discretization Points per Spatial Dimension
$d$ Physical Dimension of the TT cores
$D$ Spatial Dimension
$r$ TT rank
$\delta_{\mathrm{tol}}$ Accuracy Threshold for the Algorithms
$\delta_{\mathrm{acc}}$ Approximation Error for the TT
$\delta_{\mathrm{trunc}}$ Truncation Threshold for the SVD
$\varepsilon$ Strain Field
$\sigma$ Stress Field
$\varGamma_{0}$ Green-Eshelby Operator
$\mathbf{C}_{(0)}$ (isotropic) Stiffness Tensor
In this paper, we will combine the adaptation of Dolgovs SFFT algorithm with Tensor Network and Homogenization Theory to obtain a novel quantum-inspired SFFT-based Homogenization algorithm. We will further explore the applicability of our method to various linear elastic geometries in both 2D and 3D. A key aspect of our analysis is the memory consumption and speed of our approach, which we compare to the state-of-the-art FFT-based Homogenization method. Our goal is to achieve scalability similar to the QFT while ensuring practical computability with classical computers. Additionally, our approach has the potential to generate the large amount of training data needed for deep material networks (DMNs) [23, 24, 25]. In their original formulation, DMNs are trained exclusively on linear elastic data, yet they are successfully applied to nonlinear and inelastic problems with remarkable accuracy.
The outline of the remainder of this paper is as follows: Section 2 presents the derivation of the FFT-based Homogenization algorithm and discusses its limitations, while Section 3 introduces the Tensor Train format. Section 4 presents the ’Zip-Up’ algorithm, which converts our QFT quantum circuit into the desired SFFT Tensor Train Operator, and concludes with our novel SFFT-based Homogenization algorithm. In Section 5, we compare the performance of our approach to the state-of-the-art FFT-based method, focusing primarily on speed and memory consumption. Before closing, we provide a short summary of the paper – along with the statement of the main conclusions drawn – in Section 6. For a comprehensive list of the most important symbols and abbreviations used within this manuscript, see Table 1.
@@ 2 FFT-based Homogenization
Operations Rank Result Time Complexity
$z=x*\mathnormal{const}$ $r(z)=r(x)$ $\mathcal{O}\left(dr(x)\right)$
$z=Ax$ (solve z) $r(z)$ $\mathcal{O}\left(ndr(A)r^{3}(z)\right)$
$Az=x$ (solve z) $r(z)$ $\mathcal{O}\left(ndr(A)r^{3}(z)\right)$
We continue the derivation by identifying that the local stress tensor must satisfy the force-equilibrium equation, given by
@@
[…]
@@
Working with tensor representations can often be cumbersome, both in terms of clarity and computational efficiency. To address this, we adopt Voigt notation, which represents stress and strain tensors as 3 (6) dimensional vectors, while the stiffness tensor is represented as a $3\times 3$ ($6\times 6$) matrix that operates on them, within a 2 (3) dimensional space. This reduction is made possible by exploiting the inherent symmetries of the tensors [31].
@@
The effective stiffness tensor $\mathbf{C}^{\mathrm{eff}}$ can conveniently be determined in Voigt notation, where the strain-stress relationship (1) reduces to a simple matrix-vector product. By selectively controlling the mean strain $\mathbf{E}$ where only its $k$-th component is set to unity while all others are zero - the relationship simplifies to its component form
@@
We conclude this Section by noting that the computational complexity of the FFT-based algorithm 1 is dominated by the FFT operations, scaling with $\mathcal{O}\left(DN^{D}\log N\right)$, where $N$ is the number of discretization points per dimension $D$. Thus, improving the FFT-based Homogenization procedure relies on either enhancing the performance of FFT operations, such as through hardware acceleration, or circumventing the classical FFT entirely.
@@ 3 Tensor Train Methods
Tensor Networks (TNs) have a rich and diverse history, having been developed independently in both the mathematics and physics communities [32, 33, 34]. This cross-disciplinary development led to a variety of powerful and complementary approaches across the field. In the context of TNs, high-order tensors are decomposed into a network of lower-order tensors, interconnected through specific contraction patterns. This allows for efficient manipulation and flexible representation of complex data.
This work specifically focuses on the Tensor Train (TT) format, along with some of its special cases, for representing high-order tensors $\mathcal{T}$ [35, 36]. The elements of $\mathcal{T}$ are related to its TT decomposition as follows
@@
where each physical index $i_{k}$ runs from 00 to $d_{k}-1$. The novel 3-order tensors $C^{\alpha_{k}}_{\alpha_{k-1}}(i_{k})$ of shape $(r_{k}\times d_{k}\times r_{k-1})$ are regarded as the cores of the TT. The highest dimensionality of all the virtual subspaces is denoted as the rank $r$ of the TT, providing insight into the feasibility of employing TTs.
The primary advantage of TNs over full tensor representations lies in their data compression capabilities. An $n$-order tensor, with each index having a dimension of $d$, requires exponential memory $O(d^{n})$ to store its data. In contrast, the same tensor in the TT format consists of $n$ cores and scales linearly with both its dimension and its number of cores $O(ndr^{2})$, while now additionally depending on its rank. Thus, effective compression is achieved if the rank $r$ remains at a moderately level.
Moreover, TTs serve not only as a means of compressed storage. They offer a robust mathematical framework that enables efficient operations, leading to significant computational speedups when applicable. The extent of this acceleration depends primarily on the TT rank. The time complexity and rank evolution of TTs under basic linear algebra are summarized in Table 2. Given the broad mathematical foundation of TTs, we refer readers to I. V. Oseledets introduction for a comprehensive discussion for the basic linear algebra required [37].
A generalization called the Tensor Train Operator (TTO) can be obtained by extending the TT formula (9) to include an additional physical index per core
@@
The physical index pairs of each core can be interpreted as incoming and outgoing dimensions, which allows a TT to be contracted with a TTO. Therefore, the TTO serves as a complex transformation on a TT, and its contraction with a TT is commonly referred to as the ’matrix-vector’ product.
A more complex class of operations includes optimization algorithms, often employed to solve linear systems of equations (LSE). In this work, rather than employing such techniques for solving LSEs directly, we adapt the alternating minimal energy (AMEn) algorithm [38] to efficiently compute matrix-vector products. This adaptation modifies the cost function in the optimization process, allowing us to obtain an approximate solution $v$ of the equation $Au=\tilde{v}$, with the exact TTs $u$ and $\tilde{v}$ as well as the TTO $A$. The adaptation specifically minimizes the following cost function
@@
The general derivation closely follows the procedure outlined in the original AMEn paper [38], with deviations in relevant details. Additional information on the adapted method is provided in Appendix B.
The primary advantage of this class of optimization methods over direct contractions between a TTO and a TT becomes evident in the high-rank regime. The introduced approach not only reduces the overall time complexity but also leads to lower final ranks, often eliminating the need for an additional rounding step, which is discussed further below. As shown in Table 2, this significantly speeds up subsequent operations.
To efficiently perform linear algebra operations between TTs, converting a regular TT into a TTO can be beneficial. This is mainly due to the enhanced efficiency of the aforementioned optimization methods. Assume we have a TT $u$ as defined by equation (9). We can raise it to a TTO $U$, following the format in equation (3), by employing delta functions
@@
where the novel TTO cores $G^{\alpha_{k}}_{\alpha_{k-1}}(i_{k},j_{k})$ are 4-order tensors of shape $(d_{k}\times r_{k}\times d_{k}\times r_{k-1})$. The TTO derived from relation (11) applied to a TT $v$ produces the same outcome as the Hadamard product of $u$ and $v$. However, the rank of the resulting TT exhibits improved scalability, as it is no longer governed by the product of the ranks of $u$ and $v$, see Table 2. A detailed proof is provided in Appendix C.
In general, operations increase the rank when applied on a TT, eventually making the format computationally intractable, as shown in Table 2. To address this issue, additional rounding operations must be employed. Given a tensor $v$ in TT format, the goal is to find an approximation $\tilde{v}$ such that
@@
fulfilling the desired accuracy $\delta_{\mathrm{acc}}$ in terms of the frobenius norm ||$||_{F}$. The process involves sequential orthogonalization of TT cores followed by truncated Singular Value Decompositions (SVDs), cutting small singular values. To reach the desired accuracy, the truncation threshold during the procedure must satisfy the condition
@@
This iterative process preserves essential information while maintaining a reduced rank, thereby ensuring computational tractability [37].
@@ 4 Quantum-Inspired Approach
To align the TT format with the mathematical formulation of quantum circuits, we restrict the dimensionality of the physical indices to $d=2$. This specialized form, known as the Quantized Tensor Train format, directly links the dimensionality of the TT cores to the single-qubit subspaces within quantum circuits. Thus, an $n$-core TT represents an $n$-qubit quantum state, with unitary gates acting on these qubits being represented as Quantized Tensor Train Operators. To improve readability, we will use the abbreviation TT(O) to additionally refer to its specialized variants. However, the specific version in use will always be the one explicitly defined most recently.
The mathematical relationship between the TT format and quantum circuits reveals a closeness in both their mathematical structure and diagrammatic representation. This duality forms the cornerstone of our quantum-inspired approach. However, although the TTOs are derived from unitary operators, they are generally non-unitary. This is due to the cut-off of singular values introduced during the TT rounding step to ensure computational tractability. Further, quantum phenomena like entanglement are a special challenge for the TT format, since they result in an increased rank and thus an additional computational overhead. If entanglement grows too large, Tensor Network based methods in general are likely to reach their computational limits [39].
One major advantage of quantum-inspired approaches is that they operate entirely on classical computers. This not only avoids the high error rates of NISQ-era quantum systems but also circumvents the significant challenges associated with transferring classical information to and from quantum computers. [40].
Figure 1: Quantum Circuit for the Quantum Fourier Transform $\mathcal{QFT}_{n}=\mathcal{R}_{n}\mathcal{Q}_{n}$ for $n=4$ [41].
@@ 4.1 The Quantum Fourier Transform
[…]
@@
The complete circuit is shown in Figure 1. Alternatively, the QFT can be written as a product of two operators $\mathcal{QFT}_{n}=R_{n}\,Q_{n}$. The $Q_{n}$ circuit is responsible for the transformation itself, while the $R_{n}$ contribution reorders the output through a succession of $SWAP$ Gates. The latter one is needed to achieve the same ordering convention as is used within its classical counterpart.
The QFT exhibits an exponentially improved time complexity of $\mathcal{O}\left(D\log^{2}N\right)$, where $N$ is the number of discretization points per dimension $D$, compared to the FFT. The latter one scales according to $\mathcal{O}\left(DN^{D}\log N\right)$, making the QFT a desirable alternative.
The transfer of the QFT into its corresponding TTO formulation has been done previously, however, it was deemed impractical due to its high ranks [22, 42]. Recently, Chen et al. [21] showed that this may be true for the full QFT, but not for its reduced form $\mathcal{Q}_{n}$. Specifically, they showed that when the reordering operator $\mathcal{R}_{n}$ is not considered, the ranks obtained for the TTO formulated reduced QFT decay exponentially with $n$.
This version of the QFT, known as the Superfast Fourier Transform (SFFT) [22, 21], has been observed to significantly outperform the FFT in certain cases, once an input-dependent crossover point is reached. However, these cases require the underlying data to be (partially) structured.
We adopt a similar approach by focusing on the reduced form of the SFFT, which will be derived in the upcoming section. However, this results in a different momentum space convention, which will be addressed in the derivation of the SFFT-based Homogenization algorithm. It is worth noting that on current NISQ-era quantum hardware, the reordering operation can similarly cause issues due to the numerous error-prone SWAP operators [43].